OpenStudy (anonymous):

Add and simplify x+3/x+4 + x+5/x+4

OpenStudy (anonymous):

denominators are common so you can just add = x+3 + x+5/x+4 = 2x+8/x+4 = 2(x+4)/(x+4) = 2

OpenStudy (anonymous):

2(x+4)^2 -------- x